OpenStudy (anonymous):

1/2x + 5/2(x - 2) = -3 - 1/2x - 2(x + 1) pls help meeeeee

OpenStudy (anonymous):

x=0

jhonyy9 (jhonyy9):

x+5x-10=-6-x-x-1 8x=3 x=3/8